Given a force vector F1=0.3kg⋅9.81ms2, at 59 degrees with respect to the x axis and a second vector force F2=0.2kg⋅9.81ms2, at 18 degrees with respect to the x axis the resultant of these 2 forces has a magnitude of:


A) 2.9N;
B) 120.8N;
C) 52.0N;
D) 4.6N;
Answer #1

F, = (0-3 269.81) = 2.943 N F, = (0.2)(9.81 ) = 1.962N o = Angle between both Forces - 59 - 18 = 419 Face - VER AFP + 2 Fife